# “Would You Want an AI Tutor?” Understanding Stakeholder Perceptions of LLM-based Chatbots in the Classroom

### v1

@@ Appendix B Models of Perception
Complete breakdown of models and theories conceptualizing attitudes, beliefs, and/or behaviors toward technology, as they appear in the literature, to investigate stakeholders’ perceptions of LLM-based chatbots. The breakdown supports the thesis outlined in the third gap identified in the literature in Section 3.2.3, namely, the lack of homogeneity in the assessment of perceptions and the need for standardization in how perceptions are conceptualized.
@@
Appendix C Perception Scales
@@
Short Version of the User Experience Questionnaire (UEQ-S) The User Experience Questionnaire (UEQ) aims at collecting end-users’ feelings, impressions, and attitudes that arise when experiencing a product. It consists of 26 items grouped into 6 scales, namely Attractiveness, Perspicuity, Efficiency, Dependability, Stimulation and Novelty (schrepp2017design). Schrepp, Hinderks, and Thomaschewski later developed a Short Version of the UEQ, reducing it to 8 items grouped into two scales, which measure two meta-dimensions: pragmatic and hedonic quality, while still covering the spectrum of product qualities assessed by the original UEQ.
@@
Pupils’Attitudes Towards Technology Short Questionnaire (PATT) Pupils’Attitudes Towards Technology (PATT) was first developed by de1988techniek and has been recently reconstructed and revalidated by Ardies et al.. The revision led to the creation of the PATT Short Questionnaire (PATT-SQ), where the original 58 items were reduced to 24, divided into six subfactors. These six subfactors are Career Aspirations, Interest in Technology, Tediousness of Technology, Positive Perception of Effects of Technology, Perception of Difficulty, and Perception of Technology as a Subject for Boys or for Both Boys and Girls. The PATT-SQ was examined and further developed by Svenningsson et al., leading to the creation of the PATT-SQ-SE.

### v2

@@ Appendix C Models of Perception
Complete breakdown of models and theories conceptualizing attitudes, beliefs, and/or behaviors toward technology, as they appear in the literature, to investigate stakeholders’ perceptions of LLM-based chatbots. The breakdown supports the thesis outlined in the third gap identified in the literature, namely, the lack of homogeneity in the assessment of perceptions and the need for standardization in how perceptions are conceptualized.
@@
Appendix D Perception Scales
@@
Short Version of the User Experience Questionnaire (UEQ-S) The User Experience Questionnaire (UEQ) aims at collecting end-users’ feelings, impressions, and attitudes that arise when experiencing a product. It consists of 26 items grouped into 6 scales, namely Attractiveness, Perspicuity, Efficiency, Dependability, Stimulation and Novelty (Schrepp et al., 2017). Schrepp, Hinderks, and Thomaschewski later developed a Short Version of the UEQ, reducing it to 8 items grouped into two scales, which measure two meta-dimensions: pragmatic and hedonic quality, while still covering the spectrum of product qualities assessed by the original UEQ.
@@
Pupils’Attitudes Towards Technology Short Questionnaire (PATT) Pupils’Attitudes Towards Technology (PATT) was first developed by Vries, de and has been recently reconstructed and revalidated by Ardies et al.. The revision led to the creation of the PATT Short Questionnaire (PATT-SQ), where the original 58 items were reduced to 24, divided into six subfactors. These six subfactors are Career Aspirations, Interest in Technology, Tediousness of Technology, Positive Perception of Effects of Technology, Perception of Difficulty, and Perception of Technology as a Subject for Boys or for Both Boys and Girls. The PATT-SQ was examined and further developed by Svenningsson et al., leading to the creation of the PATT-SQ-SE.